Grade 11Physical Chemistry

energy of an electron is given by E=-2.178×10-18j (z2÷n2).wavelenght of light require to excite in an electron in an h atom from level n=1 to n=2 will be

gohilmukesh
8 Years agoGrade 11

1 Answer

IHate Ppts
8 Years ago
According to Rydberg's equation
1/λ = RZ2(1/n12-1/n22)
Substituting values,
1/λ = R(1)2(1/12 - 1/22)
1/λ =109677(1-1/4)
1/λ =82257.75
λ =1.215 X 10-5